# DaBiT: Depth and Blur informed Transformer for Joint Refocusing and Super-Resolution

### v1

@@ 4.2 DaBiT
Transformers are undoubtedly an excellent tool for video restoration, but their high memory footprint and slow inference speed poses problems in many use cases. The novel sparse transformer blocks proposed by Zhou et allet@token. [zhou2023propainter] partially overcome this issue by building on the window-based transformer blocks proposed in E2FGVI [li2022towards] and FGT [zhang2022flow]. Hence, we adopt a similar architecture to that of ProPainter, but with a few modifications. Firstly, in addition to the input frames and binary masks, we encode the depth and blur maps. This enables the transformer to learn the spatial relation between blur and depth, and ensures high similarity by encouraging consistent depth across frames. We encode this fused input into the feature space in order to further speed up processing. In contrast to ProPainter, a learnable feature propagation module is not employed before the transformer due to the minimal performance loss when ablated.
@@
We then use soft split [liu2021fuseformer] to generate patch embeddings $P$, and decompose them into $m\times n$ non-overlapping windows of dimension $h\times w$. This results in $P_{win}$, which we obtain the query, key, and value matrices from. Following [zhou2023propainter], we design separate strategies for the query and key/value spaces as described below. We also integrate global tokens [zhang2022flow] and the window expand strategy [liu2021fuseformer] in order to replicate the efficient $5\times 9$ window size.
@@
Query Space. Since we are in effect attempting to inpaint a very large mask (up to 100% of the frame), sparsity will not always result in as large a performance saving. However, we hypothesise that communicating to the model the areas which are more or less blurry, and how blurry those regions are, will aid the transformer’s self-attention mechanism. However, savings are still possible from any regions not exhibiting blur. To this end, we downsample the blur maps and apply MaxPooling across frames to obtain a query mask $M_{Q}$, which is zero at a pixel $(x,y)$ if the frames in question are all in focus at that point. The mask additionally communicates which areas of the frames are most blurred, assisting the attention mechanism by effectively acting as a form of positional encoding for each patch.
@@
Key and Value Space. We employ the same strategy as [zhou2023propainter], and incorporate a temporal stride of two, effectively dividing computation of frames evenly between each of the eight transformer blocks. This is primarily motivated by the relative similarity of defocus between two given frames, and the 50% compute savings during training and inference associated with it.
@@
Super Resolution. Finally, we add an additional upsampling module inspired by [chan2022basicvsr++] consisting of multiple cascading PixelShuffle layers to achieve 2$\times$ video super-resolution. This component allows for flexibility in input resolution without unnecessarily increasing model size or speed.
@@ 4.3 Loss functions
where $I_{pred}$ and $I_{gt}$ are the predicted frames and the ground-truths, respectively, and $\epsilon=0.001$. The second loss term denotes the $l_{1}$ loss between only the blurred areas of the input frames and the corresponding regions in the ground truth frames. The third term is the inverse of the second, comparing the focused input regions to the focused ground truth regions. These two loss terms are inspired by various inpainting approaches, and encourages the model to remove blur while remaining true to the in-focus regions. Thus, in total, the loss is expressed as
$\mathcal{L}_{tot}=\alpha\mathcal{L}_{Charb}(I_{pred},I_{gt})+\beta l_{1}(\hat{I}_{pred},\hat{I}_{gt})+\gamma l_{1}(\tilde{I}_{pred},\tilde{I}_{gt}),$ (2)
@@
where $\hat{I}$ denotes only the blurred regions of the frame sequences, and $\tilde{I}$ denotes only the focused regions. Both of two the $l_{1}$ loss term inputs are resized to the input resolution, enabling the model to be more sensitive to the numerically small values of the Charbonnier super-resolution loss term. For training, we set $\alpha=\beta=\gamma=1$.
5 Experiments
@@
Training Details and Metrics. We train our models on both the Youtube-VOS [xu2018youtube] and BVI-DVC [ma2020bvi] datasets. During the training of the Flow Refocusing module, 20 iterations of RAFT [teed2020raft] are used for flow estimation with a frame sequence length of 10. We follow [zhou2023propainter] and apply propagation to 8$\times$ downsampled feature vectors. In the main model, eight sparse transformer blocks with four heads are employed for deblurring a local video sequence $l=10$, with $r=6$ reference frames. The map guided transformer window size is 5$\times$9, and the extended size is half of the window size. We train both DaBiT and the Flow Refocusing module using the Adam optimiser [kingma2014adam] with a batch size of 1 on frames of size 432$\times$240, due to memory constraints. We fix the learning rate at $10^{-4}$, and train the Flow Refocusing module for 100k iterations and DaBiT for 300k iterations on a single NVIDIA RTX 4090 GPU. Following  [youk2024fma], we use PSNR and SSIM [wang2004image] to evaluate our methods, as well as tOF [chu2020learning] for measuring temporal consistency.
@@
Figure 4: Qualitative examples from blurred BVI-Lowlight scenes showing the performance of our approach on lowlight data. The brightness has been increased by a factor of two for columns (b)–(f) for better visibility. Superscript (*) indicates models that have been retrained to remove focal blur using our training approach.
Approach Model PSNR $\uparrow$ SSIM $\uparrow$ tOF $\downarrow$ #P RT
Video Inpainting ProPainter [zhou2023propainter] 17.962 0.597 4.992 39.4 0.107
Video SR BasicVSR++ (120p) [chan2022basicvsr++] 25.186 0.679 3.983 7.3 0.022
BasicVSR++ [chan2022basicvsr++] 25.618 0.713 3.716 7.3 0.150
Image Refocusing Restormer [zamir2022restormer] 24.715 0.698 2.902 26.5 0.093
APL [zhao2022united] 23.870 0.605 4.768 13.5 0.017
Joint Refocusing FMA-Net* [youk2024fma] 25.691 0.721 2.396 9.6 0.543
FMA-Net (120p)* [youk2024fma] 26.860 0.720 2.291 9.6 0.151
DaBiT (Ours) 28.777 0.811 1.239 45.5 0.112
Table 1: Quantitative comparison results for DaBiT and three other tested methods. For each column, the best result is bold in red and the second best is underlined in blue. The average runtime (RT) for refocusing a 240p frame as well as the number of model parameters (#P) for each method are also reported. For super-resolution models, the output is downsampled to match the 480p ground truth DAVIS scene, and for non-super-resolution models the ground truth is downsampled. Superscript (*) indicates models have been retrained using our defocus blur training approach.
5.1 Comparison
Qualitative Evaluation. Visual samples of our model’s outputs on DAVIS-Blur are shown in Figure 3. It can be observed that the DaBiT outputs are considerably more detailed than those from other methods. Additionally, our model shows promise on low-light defocus blur, as tested on select scenes from the BVI-Lowlight dataset [anantrasirichai2024bvi], shown in Figure 4. In this case, we create a similar test environment to that of DAVIS-Blur, with an initial focal point $f=0$, a maximum blur kernel of $7\times 7$, and a focal range of 100.
Quantitative Evaluation. We quantitatively compare our method against several state-of-the-art methods in Table 1. For a fair comparison, deblurring methods not originally designed for removing defocus blur are finetuned using our training approach for 35 epochs. Due to the 4$\times$ upscaling nature of FMA-Net and BasicVSR++, we run the models in two ways. In the first, we pass a 240p frame and downsample the output to match the ground truth, and in the second we pass a 120p frame for comparison to the 480p ground truth. The PSNR, SSIM, and tOF are reported for each scenario.
@@ 5.2 Ablation Study
Experiment PSNR $\uparrow$ SSIM $\uparrow$ tOF $\downarrow$
w/o Image Propagation or Blur Maps 25.227 0.703 2.548
w/o Blur Maps 25.311 0.706 2.499
w/o Image Propagation 27.203 0.758 1.914
DaBiT (Ours) 28.777 0.811 1.239
Table 2: Ablation study on the effectiveness of image propagation and the use of blur maps on DAVIS-Blur.
Blur Maps and Propagation. To demonstrate the importance of both blur maps and image propagation in our model, we conduct an ablation study with these components removed. Since the image propagation module is not learned, and simply updates the masks and blurry inputs, our model does not need to be retrained. The same is true for the way in which we ablate the blur map by replacing the variable with a tensor of all ones. This is essentially telling the model that the entirety of every frame is blurred, preventing the use of learned performance from blur maps.
As shown in Table 2, the propagation components benefit our model, but less so than the blur maps. Additionally, the model with neither image propagation nor blur maps exhibits a somewhat surprising tOF score - similar to the model without blur maps. This is likely due to the inability of the model to remove the blur present in the video, since a high degree of blur inherently improves temporal consistency.
@@ 6 Conclusion
This paper presents a novel approach to the relatively understudied problem of video focal deblurring. Our method utilises a novel transformer and flow refocusing module, along with bidirectional image propagation, to remove focus blur from a given video sequence. We overcome the limitations of computational resources and speed by working with low-resolution inputs and embedding a super-resolution module in the pipeline. Additionally, we propose a new method to generate synthetic focal blur data, enabling the model to learn from a broader range of content. When compared to other joint video deblurring and super-resolution models on the proposed DAVIS-Blur test set, state-of-the-art performance is achieved. In addition, our model exhibits excellent performance on low-light footage, which often encounters focusing issues. Future work will focus on generalising this approach to other forms of blur, increasing the super-resolution scaling, and investigating if similar performance can be achieved at increased speed by a smaller model using model compression.
7 Acknowledgements
This work was jointly funded by the UK EPSRC (iCASE Award), the BBC, the UKRI MyWorld Strength in Places Programme, and the University of Bristol.
[…]
Appendix A Appendix 1
Figure 5: The blurring procedure adopted during training.
Figure 5 shows the blurring process we employ during the training of our model. A high resolution input sequence of length $l$ is selected along with the corresponding depth maps. A random focal point $f$ and focal range $f_{r}$ are chosen, as well as a random focus rate $\frac{df}{dt}$ and maximum blur size $n_{max}$. These parameters define a focus shift’s size, intensity, and speed, and the focus shifts in the direction it has furthest to travel; in this case, away from the camera. The unfocused area is blurred according to its distance from the focal range; nearer areas are blurred less, and the furthest distances are blurred at the maximum blur size, $n_{max}$. Note that the rectangular blur regions are an approximation, and the actual regions of the blur at each depth are shown in the generated ground-truth depth map. Finally, we downsample the blurred sequence, along with its blur and depth maps, to obtain a blurred input for training.

### v3

@@ 3.4 Map Guided Transformer
Transformers are undoubtedly an excellent tool for video restoration, but their high memory footprint and slow inference speed pose problems in many use cases. The novel sparse transformer blocks proposed by Zhou et al. [56] partially overcome this issue by building on the window-based transformer blocks proposed in E2FGVI [27] and FGT [53]. Hence, we adopt a similar architecture to that of ProPainter, but with a few modifications. Firstly, in addition to the input frames and binary masks, we encode the depth and blur maps. This enables the transformer to learn the spatial relation between blur and depth, and ensures high similarity by encouraging consistent depth across frames. We encode this fused input into the feature space in order to further speed up processing. In contrast to ProPainter, a learnable feature propagation module is not employed before the transformer due to the minimal performance loss when ablated.
Figure 3: Map guided transformer. To decrease memory consumption and increase speed, sparsity is introduced in the key, values, and query spaces. The blur maps inform the attention mechanism of the spatio-temporal behavior of focal blur.
@@
Soft split [32] is used to generate patch embeddings $P$ in the feature space, and decompose them into $m\times n$ non-overlapping windows of dimension $h\times w$. This results in the set of patches $P_{win}$, which we obtain the transformer’s query ($Q$), key ($K$), and value ($V$) vectors from. Following [56], we design separate sparsity strategies for the query and key/value spaces as described below. We also integrate global tokens [53] and the window expand strategy [32] in order to replicate the $5\times 9$ window size, the smallest possible, due to the speed achieved by  [56]. An overview of this process is shown in Figure 3.
@@
Query Space. We hypothesize that communicating to the model the areas which exhibit more blur or less blur, and precisely how blurry those regions are, will aid the transformer’s performance. However, compute savings are possible from any regions not exhibiting blur, since they do not need to be refocused. To this end, we downsample the blur maps and apply MaxPooling across the temporal dimension to obtain a query mask $M_{Q}$, which is zero at a pixel $(x,y)$ if the frame in question is correctly focused at that point. Crucially, the continuous mask communicates which areas of the frames are most blurred, thereby assisting the attention mechanism in its encoding of each patch.
@@
Key and Value Space. The key and value space incorporates a temporal stride of two, following the same strategy as [56], effectively dividing computation evenly between each of the eight transformer blocks. This is primarily motivated by the relative similarity of blur between two given frames, and the 50% compute savings during training and inference.
@@
Super Resolution. Finally, we add a simple upsampling module consisting of three cascading PixelShuffle layers to achieve 2$\times$ video super-resolution. This component allows for flexibility in input resolution without unnecessarily increasing model size or speed.
@@ 3.5 Loss functions
where $I_{pred}$ and $I_{gt}$ are the predicted frames and the ground-truths, respectively, and $\epsilon=0.001$. The second loss term applies an $l_{1}$ loss specifically to the blurred areas, encouraging the model to prioritize restoring these regions while leaving focused areas intact. The third term is the inverse of the second, comparing the focused input regions to the focused ground truth regions. These two loss terms are inspired by various inpainting approaches, and encourage the model to remove blur while remaining true to the in-focus regions. Thus, in total, the loss is expressed as
$\begin{split}\mathcal{L}_{tot}&=\mathcal{L}_{Charb}(I_{pred},I_{gt})\\ &+l_{1}(\hat{I}_{pred},\hat{I}_{gt})+l_{1}(\tilde{I}_{pred},\tilde{I}_{gt}),\end{split}$ (2)
@@
where $\hat{I}$ denotes only the blurred regions of the frame sequences, and $\tilde{I}$ denotes only the focused regions. Both of the two $l_{1}$ loss term inputs are resized to the input resolution, enabling the model to be more sensitive to the numerically small values of the Charbonnier super-resolution loss term.
4 Training Content Generation
Deep learning necessitates a substantial volume of data to train the model effectively. While unsupervised learning techniques exist, supervised learning-based methods still outperform them in terms of both accuracy and speed. However, obtaining paired datasets of blurred and clean sequences is time-consuming and content diversity can be limited. Hence, we introduce a new approach to generate diverse synthetic training material for optimizing the proposed method.
4.1 Focal Blur Simulation
In order for our model to learn the nature of focal blur, an accurate blur model is required. We begin with an input video $I_{gt}=\{I_{t}\in\mathbb{R}^{3\times H\times W}\}_{t=1}^{T}$ with $T$ frames. We then compute the depth maps $D_{gt}=\{D_{t}\in\mathbb{R}^{1\times H\times W}\}_{t=1}^{T}$ for the input using DepthAnything [48], due to its ability to predict temporally consistent depths at a level surpassing other state-of-the-art models such as MiDaS [37] or ZoeDepth [3]. Using the depth maps, the spatial variance of focal blur is able to be modeled. We begin by setting the focal point at a depth $f$, and defining a focal range $f_{r}$ centered on the focal point. We then apply gaussian blur kernels $G(x,y)=\frac{1}{2\pi\sigma^{2}}e^{-\frac{x^{2}+y^{2}}{2\sigma^{2}}}$, with $\sigma=5$, to the image in both directions along its depth axis, scaling symmetrically from zero inside the focal range to a maximum kernel size of $n_{max}\times n_{max}$. In this way, we blur the furthest depths from the focal point the most, creating a realistic focal blur effect.
During training, a random frame sequence from $I_{gt}$ of length $l,l<T$, is taken, along with $r$ random global reference frames from elsewhere in the video. A random focal point ($f$) and focal range ($f_{r}$) are selected, along with a random focus rate $\frac{df}{dt}$, dictating the speed and direction with which the focal point will change in the sequence. A random maximum blur kernel size $n_{max}\in\{x\in\mathbb{Z}:3\leq x\leq 11,x\text{ is odd}\}$ is chosen, and the sequence and its global reference frames are blurred in a temporally variant manner, simulating a focus shift. The spatial blur intensity information is recorded and used to create a ground truth blur map $B_{gt}=\{B_{t}\in\mathbb{R}^{1\times H\times W}\}_{t=1}^{T}$, assisting our model during training. We also introduce data augmentation by means of a random horizontal flip and random reversal of sequence during the training. Finally, to enable super-resolution, we downsample the blurred sequence, along with its blur maps, denoted as $\widetilde{I_{in}}$ and $B_{in}$, respectively. This approach is illustrated by Figure 4.
Figure 4: The training content generation process.
The variation in blur intensity and focus speed simulates a broad range of focus shifts. Some are representative of sequences which do not contain any additional global information (e.g. $\frac{df}{dt}=0$, so $l$ is out-of-focus in the same spatio-temporal regions), and some which contain enough focused footage to be able to nearly reconstruct the whole sequence using only image propagation (e.g. $f$ is varied completely across $\text{min}(D_{gt})$ and $\text{max}(D_{gt})$). This allows a given model to learn to deblur fully out-of-focus frames while not discarding the information contained in focused frames.
Approach Model PSNR (dB) $\uparrow$ SSIM $\uparrow$ tOF $\downarrow$ #P (M) RT (s)
Video Completion ProPainter [56] 17.962 0.597 4.992 39.4 0.107
Video SR BasicVSR++ (120p) [5] 25.186 0.679 3.983 7.3 0.022
BasicVSR++ [5] 25.618 0.713 3.716 7.3 0.150
Image Refocusing Restormer [50] 24.715 0.698 2.902 26.5 0.093
APL [54] 23.870 0.605 4.768 13.5 0.017
Joint Refocusing FMA-Net* [49] 25.691 0.721 2.396 9.6 0.543
FMA-Net (120p)* [49] 26.860 0.720 2.291 9.6 0.151
DaBiT (Ours) 28.777 0.811 1.239 45.5 0.112
V1 w/o Blur Maps 25.311 0.706 2.499
V2 w/o Image Propagation 27.203 0.758 1.914
V3 w/o Blur Maps and Image Propagation 25.227 0.703 2.548
Table 1: Quantitative comparison results for DaBiT (and its ablation study variants) and three other tested methods. For each column, the best result is bold in red and the second best is underlined in blue. The average runtime in seconds (RT) for refocusing a 240p frame as well as the number of model parameters in millions (#P) for each method are also reported. For super-resolution models, the output is downsampled to match the 480p ground truth DAVIS scene, and for non-super-resolution models the ground truth is downsampled. Superscript (*) indicates models have been retrained using our defocal blur training approach.
To generate the training material, we combined the 3471 sequences from the Youtube-VOS [45] and the 200 sequences of the BVI-DVC [33] datasets. Youtube-VOS is a large video database designed for object segmentation, while BVI-DVC contains diverse content with a wide range of low-level features and semantic scenarios, which has been employed to optimize learning-based models for various low-level tasks [12, 34, 52]. This results in a total of 3671 unique videos.
5 Experimental setup
@@
Training Details and Metrics. During the training of the Flow Refocusing module, 20 iterations of RAFT [41] are used for flow estimation with a frame sequence length ($l$) of 10. We follow [56] and apply propagation to 8$\times$ downsampled feature vectors. In the main model, eight sparse transformer blocks with four heads are employed for deblurring a local video sequence $l=10$, with $r=6$ reference frames. The map guided transformer window size is 5$\times$9, and the extended size is half of the window size.
Both DaBiT and the Flow Refocusing module are trained using the Adam [22] optimizer, selected for its robust performance in optimizing deep learning models. A fixed learning rate of $10^{-4}$ and a batch size of 1 are used for both models to ensure gradual convergence. Due to memory constraints, we train the Flow Refocusing module and DaBiT on frames of size 432$\times$240, and continue for 100k iterations and 300k iterations, respectively Benchmark methods. We benchmark DaBiT against a variety of state-of-the-art image and video restoration models, including video inpainting and completion model ProPainter [56], video super resolution approach BasicVSR++ [5], two image refocusing methods, Restormer [50] and APL [54], and one joint refocusing model, FMA-Net [49].
Test datasets and metrics. In order to create suitable test content, we apply the same content generation workflow to 50 sequences in the DAVIS [36] database, which results in a new test set for refocusing, DAVIS-Blur. Instead of the randomized parameters used during training, fixed parameters are used to standardize the test set. We set $l=T$, $n_{max}=11$, $f_{r}=100$, fix the initial focal point $f=0$, and configure the focus rate to be $\text{max}(D)/l$, according to the length of each sequence. This simulates a complete focus shift from near to far, across the full length of the sequence, in all 50 sequences. It is noted that the large range of sequence lengths in the DAVIS set allows a wide range of focus shifts to be synthesized, thus posing a challenging task. Additionally, DaBiT and benchmark methods were tested on real low-light scenes, which often face focusing problems due to high noise and low contrast. The BVI-RLV [30] dataset was used to evaluate their performance in low-light scenarios.
We use PSNR and SSIM [44] to evaluate our methods, as well as tOF [9] for measuring temporal consistency. The number of parameters in the model, expressed in millions, is reported along with the average runtime for producing a single 240p frame in seconds. The complexity figures are calculated on a PC with an Intel Core i7-13700 CPU, 64GB RAM, and a single NVIDIA RTX 4090 GPU.
@@
Figure 6: Qualitative examples from blurred BVI-RLV scenes showing the performance of our approach on lowlight data. The brightness has been increased by a factor of two for columns (b)–(f) for increased visibility. Superscript (*) indicates models that have been retrained to remove focal blur using our training approach.
6 Results and Discussion
6.1 Quantitative results
Table 1 summarizes the comparison results between our method and other five state-of-the-art methods on the DAVIS-Blur dataset. For a fair comparison, deblurring methods not originally designed for removing defocal blur are finetuned on our data with our training approach for 35 epochs, while FMA-Net is fully retrained. Due to the 4$\times$ upscaling nature of FMA-Net and BasicVSR++, we run the models in two ways. In the first, we pass a 240p frame and downsample the output to match the ground truth, and in the second we pass a 120p frame for comparison to the 480p ground truth. Also for fairness, the APL [54] codebase is modified to include a significant speedup, and ProPainter is run in video completion mode.
DaBiT outperforms all other tested methods by a minimum of 1.9 dB PSNR and 0.09 SSIM. In addition, it shows very high temporal consistency across frames, with a minimum decrease of 1.05 tOF from the next best model. For our method, the time taken for estimating depth is not taken into account, but the RAFT estimate time is included, making the model competitively fast, given its size.
6.2 Qualitative results
Visual samples of our model’s outputs on DAVIS-Blur are shown in Figure 5. It can be observed that the DaBiT outputs are considerably more detailed than those from other methods. Additionally, our model shows promise on low-light focal blur, as tested on select scenes from the BVI-RLV dataset [30], shown in Figure 6. In this case, we create a similar test environment to that of DAVIS-Blur, with an initial focal point $f=0$, a maximum blur kernel of $7\times 7$, and a focal range of 100.
@@ 6.3 Ablation study
V1 w/o Blur Maps. One of the novelties in this work is the use of blur maps in the proposed DaBiT model. To test its importance, we have created a variant (V1) by replacing the variable with a tensor of all ones. This is essentially telling the model that the entirety of every frame is blurred, preventing the use of learned performance from blur maps.
V2 w/o Image Propagation. To verify the effectiveness of the image propagation component, we have removed the image and mask propagation present in the model, obtaining V2. Since the image propagation module is not learned, and simply updates the masks and blurry inputs, this version of our model does not need to be retrained.
V3 w/o Blur Maps and Image Propagation. Another variant (V3) was created by removing both blur maps and image propagation. The ablation study results are also presented in Table 1, where all three variants show inferior PSNR, SSIM and tOF performance compared to the original DaBiT model. This demonstrates the contribution of these components. Moreover, it has been observed that without blur maps (V1), the model results in a more significant performance loss (more than 3.4dB in PSNR) compared to that without image propagation (with a loss of 1.5dB in PSNR).
6.4 Limitations
While the proposed DaBiT model demonstrates impressive performance in video refocusing tasks, there are still several limitations that need to be addressed.
Blur map estimation. We found that some results contain visible artifacts in homogeneous areas. This may be due to bandlimiting, which leads to inaccurate blur map estimation. An example is shown in Figure 7. While other areas are sharpened with significantly enhanced details by our refocusing method, the sky exhibits some artifacts. However, this can be easily corrected by applying a low-pass filter on homogeneous areas.
Complex textures. In addition, the model occasionally struggles with fine-grained details such as the text present in the upper right of Figure 7. This limitation is likely due to limited text in the training data, as well as the model’s reliance on blur maps and image propagation components, which might not effectively handle all complex textures.
Computational complexity. DaBiT is a high-performance, but relatively large, model with competitive inference speed. This allows for use in time-sensitive scenarios; however, the large size imposes high memory requirements and restricts deployment of the model on low-end hardware. Future work will aim to combat these drawbacks with model compression techniques such as model pruning optimizers like OBProx-SG [7] and knowledge distillation [16, 35].
@@ 7 Conclusion
Figure 7: Refocusing result of ‘Car’ sequence, in which the sky contains visible artifacts while other areas are enhanced with more spatial details.
This paper presents a novel approach to the relatively understudied problem of video focal deblurring. Our method utilizes a novel transformer and flow refocusing module, along with bidirectional image propagation, to remove focal blur from a given video sequence. We overcome the limitations of computational resources and speed by working with low-resolution inputs and embedding a super-resolution module in the pipeline. Additionally, we propose a new method to generate synthetic focal blur data, enabling the model to learn from a broader range of content. Compared to state-of-the-art video deblurring and super-resolution models on the proposed DAVIS-Blur test set, our method achieves superior results, particularly in detail preservation and enhancing temporal consistency. The model is also highly adaptable to real-world low-light footage, a common source of focusing issues. Future work will focus on generalizing this approach to other forms of blur, increasing the super-resolution scaling, and investigating if similar performance can be achieved at increased speed by a smaller model using model compression and knowledge distillation techniques.